OpenStudy (shubhamsrg):

A building with ten storeys, each storey of height 3 metres, stands on one side of a wide street. From a point on the other side of the street directly opposite to the building, it is observed that the three uppermost storeys together subtend an angle equal to that subtended by the two lowest storeys. What is the width of the street ??

OpenStudy (dumbcow):

 choose 3 angles , A,B,C such that A-B =C \[\tan A = \frac{30}{x}\] \[\tan B = \frac{21}{x}\] \[\tan C = \frac{6}{x}\] \[\rightarrow \tan (A-B) = \tan C\] \[\rightarrow \frac{\tan A -\tan B}{1+\tan A \tan B} = \tan C\] substitute in ratios and solve for x x = 35.5